OpenStudy (anonymous):

FACTOR 5x^3+17x^2-84x-36

OpenStudy (tkhunny):

There MUST be a Positive Rational zero for this expression. 1) Do you know why? 2) Can you find it easily if it is Rational?

OpenStudy (anonymous):

okay well actually the question was 5x^4+12x^3-101x^2+48+36, and I had to solve it, finding the zeros, I found one of them which was x-1 and I divided that by the question to find the others, but I got 5x^3+17x^2-84x-36, and idk how to factor it

OpenStudy (anonymous):

1) No , 2) what does rational mean

OpenStudy (tkhunny):

1) You just count sign changes on coefficients. This one starts out positive and changes to negative. That's one sign change. There MUST be a REAL and POSITIVE zero. 2) You must know what a RATIONAL number is. Ratios of integers? 3) How did you find that (x-1) is a factor?

OpenStudy (anonymous):

1) Ohh okat 2) Oh wait yes I do, I looked through my notes 3) and I subed in one of the factors of 36

OpenStudy (anonymous):

I subbedd in 1

OpenStudy (anonymous):

then divided the whole thing by x-1

OpenStudy (tkhunny):

3) Excellent, however, you only got lucky. The possible rational zeros are ALL the factors fo 36 divided by ALL the factors of 5. Anyway, it's always a good idea to try the easy ones first. Try x = 3.

OpenStudy (anonymous):

okay got 0

OpenStudy (anonymous):

now what do I do to find the rest?

OpenStudy (tkhunny):

Reduce it again. Then we can continue out search for Rational Zeros or we can realize that it's just Quadratic and we can use the Quadratic Formula or Completing the Square to beat the other two out of it.

OpenStudy (anonymous):

reduce it by dividing it by x-3 ?

OpenStudy (tkhunny):

Right. Did you use synthetic division to evaluate it? To find that zero on the end?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got 5x^3+25x^2-26x-28 R -48

OpenStudy (tkhunny):

Oh, then you already did the division. Do you know this? Those numbers across the bottom are the coefficients you need. You can't get a remainder. You just told me the remainder is zero. How did you get a different result?

OpenStudy (anonymous):

oh yeah, hm maybe I made a mistake, let me re-do it

OpenStudy (tkhunny):

Don't redo it. Where is the table where you manged zero? That is the one you need.

OpenStudy (anonymous):

okay I did make a mistake. now I got 5x^3+27x^2-20x-12

OpenStudy (anonymous):

wait im lost, which table? We never did a table in class for this type of questions :o

OpenStudy (tkhunny):

?? Can't be. It should be only quadratic. Start with \(5x^{4}+12x^{3}-101x^{2}+48+36\) Divide the first one. \(\dfrac{5x^{4}+12x^{3}-101x^{2}+48+36}{x-1} = 5x^{3}+17x^{2}-84x-36\) Divide the second one we found. \(\dfrac{5x^{3}+17x^{2}-84x-36}{x-3} = 5x^{2}+32x+12\) Don't go back to the beginning and start over. Work with the new, reduced polynomial.

OpenStudy (tkhunny):

You should have a table if you used "Synthetic Division".

OpenStudy (anonymous):

Ohh like that, yeah I did that ! For the zeros right?

OpenStudy (anonymous):

dont we call that quotient form ?

OpenStudy (tkhunny):

Sure, you can call it that. You can call it "Steve" if you like. Just don't go back to the start.

OpenStudy (tkhunny):

How shall we find the last two zeros?

OpenStudy (anonymous):

um, not sure